A farmer receives a fertilizer recommendation suggesting an addition of 200 #N/ac. The farmer will apply this recommendation to 17.5 acres using ammonium sulfate (21% N). How many pounds of urea will the farmer apply?
Answer #1

Answer: 16667lbs of Urea.

The farmer received the recommendation to apply 200lbs of Nitrogen per are.

For 17.5 acres, the required amount would be, 17.5 × 200= 3500lbs.

Now, 21 lbs of nitrogen are present in 100lbs of Urea.

1lb of nitrogen will be present in 100/21lbs of Urea.

3500lbs of N will be present in= 100*3500/21= 16667lbs of urea.